A pair of fair dice is rolled. Let X be the random variable representing the sum of the numbers that appear.







a. Construct the probability distribution of X for a pair of dice.







b. Find P (X≥ 8).







c. Find P (X≤ 7).







d. Find the probability that X takes an even value.







e. Find P (3 ≤X≤ 10)

1
Expert's answer
2022-02-21T11:11:49-0500

There are "6^2=36" outcomes


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 1+1 & 1+2 & 1+3 & 1+4 & 1+5 & 1+6 \\
 \hdashline
 2 & 2+1 & 2+2 & 2+3 & 2+4 & 2+5 & 2+6 \\
 \hdashline
 3 & 3+1 & 3+2 & 3+3 & 3+4 & 3+5 & 3+6 \\
 \hdashline
 4 & 4+1 & 4+2 & 4+3 & 4+4 & 4+5 & 4+6 \\
 \hdashline
 5 & 5+1 & 5+2 & 5+3 & 5+4 & 5+5 & 5+6 \\
 \hdashline
 6 & 6+1 & 6+2 & 6+3 & 6+4 & 6+5 & 6+6 \\
 \hdashline
\end{array}"

a.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & x & f & P(X=x) \\ \hline
 & 2 & 1 & 1/36 \\
 \hdashline
 & 3 & 2 & 2/36 \\
 \hdashline
 & 4 & 3 & 3/36 \\
 \hdashline
 & 5 & 4 & 4/36 \\
 \hdashline
 & 6 & 5 & 5/36 \\
 \hdashline
 & 7 & 6 & 6/36 \\
 \hdashline
 & 8 & 5 & 5/36 \\
 \hdashline
 & 9 & 4 & 4/36 \\
 \hdashline
 & 10 & 3 & 3/36 \\
 \hdashline
 & 11 & 2 & 2/36 \\
 \hdashline
 & 12 & 1 & 1/36 \\
 \hdashline
Sum= & & 36 & 1 \\
 \hdashline
\end{array}"

Construct the probability distribution of X for a pair of dice.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & p(x)\\ \hline
 2 & 1/36 \\
 3 & 1/18 \\
 4 & 1/12 \\
 5 & 1/9 \\
 6 & 5/36 \\
 7 & 1/6 \\
 8 & 5/36 \\
 9 & 1/9 \\
 10 & 1/12 \\
 11 & 1/18 \\
 12 & 1/36 \\
\end{array}"

b.


"P(X\ge 8)=P(X=8)+P(X=9)+P(X=10)"

"+P(X=11)+P(X=12)"

"=\dfrac{5}{36}+\dfrac{4}{36}+\dfrac{3}{36}+\dfrac{2}{36}+\dfrac{1}{36}=\dfrac{5}{12}"

c.


"P(X\le 7)=P(X=2)+P(X=3)+P(X=4)"

"+P(X=5)+P(X=6)+P(X=7)"

"=\dfrac{1}{36}+\dfrac{2}{36}+\dfrac{3}{36}+\dfrac{4}{36}+\dfrac{5}{36}+\dfrac{6}{36}=\dfrac{7}{12}"


d.


"P(X\ is\ even)=P(X=2)+P(X=4)"

"+P(X=6)+P(X=8)+P(X=10)"

"+P(X=12)=\dfrac{1}{36}+\dfrac{3}{36}+\dfrac{5}{36}+\dfrac{5}{36}+\dfrac{3}{36}"

"+\dfrac{1}{36}=\dfrac{1}{2}"

e.


"P(3\le X\le 10)=P(X=3)+P(X=4)"

"+P(X=5)+P(X=6)+P(X=7)"

"+P(X=8)+P(X=9)+P(X=10)"

"=\dfrac{2}{36}+\dfrac{3}{36}+\dfrac{4}{36}+\dfrac{5}{36}+\dfrac{6}{36}+\dfrac{5}{36}+\dfrac{4}{36}+\dfrac{3}{36}"


"=\dfrac{8}{9}"